AC Ltd Sales ($10 per unit)          &nbsp

                            AC Ltd

Sales ($10 per unit)             160,000

Variable Costs                     112,000

Contribution Margin           48,000

Fixed Costs                  (70,000)

Net Loss               (22,000)

  1. If AC’s advertising increased by $9,000 and they wanted to achieve a Profit of $5,000, how much (in $) would their sales have to increase?
  2. What would their Profit (or Loss) be if sales increased by 25% and fixed costs were reduced by 10%?

(Show all workings)
Answer #1

Answer

1 current units sold = 160,000 / 10 = 16,000 units

Variable cost per unit = 112,000 / 16,000 = $7

Selling price per unit = $10

Advertising increase by $9,000, advertising is a fixed cost

Therefore

Total fixed cost = 70,000 + 9,000 = $79,000

Desired profit = $5,000

Sales to earn desired profit = (fixed cost + desired profit) / contribution margin ratio

Contribution margin ratio

= (48,000 / 160,000) × 100 = 30%

Sales to earn desired profit of $5,000

= (79,000 + 5,000) / 30% = $280,000

The company need a sale of $280,000 to get a desired profit of $5,000

Vompany need to increase the sales by $120,000 (from $160,000 to $280,000) to earn a desired profit of $5,000.

2 Sales increased by 25%

Sales = 16,000 + 25% = 20,000 units

Selling price = $10

Variable cost = $3

Fixed cost reduced by 10%

Fixed cost = 70,000 - 10% = $63,000

Income statement (variable costing)

Particulars $
Sales (20,000 × 10) 200,000
Less variable cost (20,000 × 7) 140,000
Contribution margin 60,000
Fixed cost 63,000
Net loss 3,000

If the sales increased by 25% and fixed cost decreased by 10% , the company operating loss is $3,000.

If YOU are like the answer please upvote.